### v2

# Poisoning Prevention in Federated Learning and Differential Privacy via Stateful Proofs of Execution

Abstract
The rise in IoT-driven distributed data analytics, coupled with increasing privacy concerns, has led to a demand for effective privacy-preserving and federated data collection/model training mechanisms. In response, approaches such as Federated Learning (FL) and Local Differential Privacy (LDP) have been proposed and attracted much attention over the past few years. However, they still share the common limitation of being vulnerable to poisoning attacks wherein adversaries compromising edge devices feed forged (a.k.a. “poisoned”) data to aggregation back-ends, undermining the integrity of FL/LDP results.
In this work, we propose a system-level approach to remedy this issue based on a novel security notion of Proofs of Stateful Execution ($\sf{PoSX}$) for IoT/embedded devices’ software. To realize the $\sf{PoSX}$ concept, we design $\sf{SLAPP}$: a System-Level Approach for Poisoning Prevention. $\sf{SLAPP}$ leverages commodity security features of embedded devices – in particular ARM TrustZone-M security extensions – to verifiably bind raw sensed data to their correct usage as part of FL/LDP edge device routines. As a consequence, it offers robust security guarantees against poisoning. Our evaluation, based on real-world prototypes featuring multiple cryptographic primitives and data collection schemes, showcases $\sf{SLAPP}$’s security and low overhead.
I Introduction
With the rise of IoT and distributed big data analytics, data produced by edge devices have become increasingly important to understand users’ behaviors, enhance the user experience, and improve the quality of service. At the same time, privacy concerns have scaled significantly fueled by the collection (or leakage) of sensitive user data [24, 31, 58, 25, 28]. To reconcile privacy and utility, several mechanisms have been proposed to enable efficient and privacy-preserving collection of data produced by (typically resource-constrained) edge IoT devices. For example, Google has proposed Federated Learning (FL) aiming to collect and train models based on user data in a distributed, lightweight, and privacy-preserving fashion [30]; Microsoft is using Local Differential Privacy (LDP) to collect statistics of sensitive telemetry data [19].
As illustrated in Fig. 1, a crucial security obstacle in the adoption of these schemes is an adversary ($\sf{\mathcal{A}dv}$) that feeds back-end aggregators with forged data, sabotaging the entire collection process even when the other participants (other edge devices and the back-end) are honest. These attacks, known as poisoning attacks, have been widely recognized and are currently considered a major problem for both FL [53] and LDP [7]. Existing mitigations are either data-driven (e.g., in the case of data poisoning detection [57, 23, 6]) or algorithmic, e.g., by making FL and LDP mechanisms more resilient against these attacks [50, 32, 55, 33]. In both cases, security is by design best-effort, since these approaches cannot detect/prevent data poisoning at its source, i.e., at the edge devices themselves (typically, resource-constrained IoT devices).
Figure 1: Poisoning in FL/LDP-based systems
Verifier ($\sf{\mathcal{V}rf}$) Prover ($\sf{\mathcal{P}rv}$) 1) Request to run $\sf{\mathcal{F}}$ 2) ${\sf{\mathcal{O}}}\leftarrow{\sf{\mathcal{F}}}()$ 3) Compute $\sigma$ 4) Respond with ($\sigma$, $\sf{\mathcal{O}}$) 5) Verify($\sigma$, $\sf{\mathcal{O}}$)
(a) $\sf{PoX}$ protocol steps
Verifier ($\sf{\mathcal{V}rf}$) Prover ($\sf{\mathcal{P}rv}$) 1) Request to run ($\sf{\mathcal{F}}$$\sf{\mathcal{I}}$) 2) ${\sf{\mathcal{O}}}\leftarrow{\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ 3) Compute $\sigma$ 4) Respond with ($\sigma$, $\sf{\mathcal{O}}$) 5) Verify($\sigma$, $\sf{\mathcal{O}}$) no $\sf{\mathcal{S}}$ leakage
(b) $\sf{PoSX}$ protocol steps
Figure 2: Standard $\sf{PoX}$ vs $\sf{PoSX}$
In this work, we propose a systematic and practical treatment to address data poisoning at its source by leveraging architectural security features of contemporary embedded systems. This ensures that raw sensed data is correctly linked with its respective local processing, ultimately generating local aggregated data whose integrity and authenticity can be verified by back-ends in LDP/FL applications.
Specifically, we build upon the recently introduced concept of Proofs of Execution (PoX) [13] for simple embedded systems. $\sf{PoX}$ allows a low-end embedded device – Prover ($\sf{\mathcal{P}rv}$) – to convince a remote Verifier ($\sf{\mathcal{V}rf}$) that a specific function $\sf{\mathcal{F}}$ has been executed successfully (i.e., from its first to its last instruction) on $\sf{\mathcal{P}rv}$. Furthermore, PoX binds obtained results (or outputs) to a timely instance of this execution. A similar notion [34] has also been explored in high-end systems (e.g., general-purpose computers and servers, as opposed to embedded devices) based on trusted platform modules (TPMs) [54].
Intuitively, $\sf{PoX}$ can convey if the result received by an FL/LDP back-end (local aggregate data) truly originates from an edge device that has obtained this data through the correct execution of the expected software, thus thwarting poisoning attacks. On the other hand, as detailed next, the adoption of $\sf{PoX}$ in FL/LDP introduces unique non-trivial challenges that require re-thinking and re-designing existing PoX methods for this purpose.
I-A On the Insufficiency of Classic $\sf{PoX}$ to Avert Poisoning
As shown in Fig. 3, PoX is a challenge-response protocol composed of the following steps:
  1. 1.
    $\sf{\mathcal{V}rf}$ sends an authenticated request containing a cryptographic challenge ($\sf{\mathcal{C}hal}$) and asking $\sf{\mathcal{P}rv}$ to execute $\sf{\mathcal{F}}$.
  2. 2.
    $\sf{\mathcal{P}rv}$ authenticates the request, executes ${\sf{\mathcal{F}}}$ obtaining output $\sf{\mathcal{O}}$, and generates a cryptographic proof $\sigma$ of this execution by measuring (signing or a MAC-ing) $\sf{\mathcal{F}}$’s implementation in program memory along with received $\sf{\mathcal{C}hal}$, produced $\sf{\mathcal{O}}$, and execution metadata that conveys to $\sf{\mathcal{V}rf}$ if $\sf{\mathcal{F}}$ execution was performed correctly.
  3. 3.
    $\sf{\mathcal{P}rv}$ returns the output and proof ($\sigma$, $\sf{\mathcal{O}}$) to $\sf{\mathcal{V}rf}$.
  4. 4.
    $\sf{\mathcal{V}rf}$ verifies whether $\sigma$ corresponds to the expected $\sf{\mathcal{F}}$ code, received output $\sf{\mathcal{O}}$, and expected execution metadata. If so, it concludes that $\sf{\mathcal{F}}$ has executed successfully on $\sf{\mathcal{P}rv}$ with result $\sf{\mathcal{O}}$.
Step 2 above must be securely implemented by a root of trust (RoT) within $\sf{\mathcal{P}rv}$ to ensure (1) temporal consistency between $\sf{\mathcal{F}}$’s measurement and its execution, (2) correctness of $\sf{\mathcal{F}}$ execution and generated $\sf{\mathcal{O}}$ at run-time, and (3) confidentiality of the cryptographic secret used to compute $\sigma$. This RoT implementation must be unmodifiable, even when $\sf{\mathcal{P}rv}$’s application software is fully compromised. The latter is typically obtained through hardware support, e.g., from ARM TrustZone [46] (see Section II-B) or similar mechanisms.
We observe that the aforementioned $\sf{PoX}$ notion has important practical limitations. It assumes $\sf{\mathcal{F}}$ to be: (1) inputless, i.e., $\sf{\mathcal{F}}$ cannot depend on inputs external to $\sf{\mathcal{P}rv}$, and (2) stateless, i.e., $\sf{\mathcal{F}}$ must not depend on states produced by prior $\sf{PoX}$ instances in $\sf{\mathcal{P}rv}$. As a result, PoX is only suitable for simple self-contained programs that may process locally collected data (via $\sf{\mathcal{P}rv}$ local I/O interfaces) but do not depend on external inputs or prior execution states. This assumption becomes problematic when attempting to apply PoX to FL/LDP.
For FL integrity, $\sf{\mathcal{P}rv}$ should prove that a training function $\sf{\mathcal{F}}$ was executed on local training dataset $D$ using $\sf{\mathcal{V}rf}$-supplied global weights $W$ and training parameters. Moreover, no portion of $D$ should be revealed to $\sf{\mathcal{V}rf}$, requiring multiple $\sf{PoX}$ instances (e.g., multiple sensing routines executed over time) to correctly produce all data points in $D$. As detailed in Section II-A, similar requirements exist in LDP. Therefore, standard $\sf{PoX}$ can not be applied in these settings.
I-B Our Contributions
To address the aforementioned limitations, this work introduces the security notion of Proof of Stateful Execution ($\sf{PoSX}$) to enable input validation and state preservation, in addition to classic PoX guarantees. The former relaxes the constraint of inputless functions in traditional $\sf{PoX}$, while the latter ensures that $\sf{PoSX}$ can use $\sf{\mathcal{P}rv}$ pre-existing states as long as they originate from a prior authentic $\sf{PoSX}$ execution. In essence, PoSX offers assurance that execution of $\sf{\mathcal{F}}$, computed with authentic input $\sf{\mathcal{I}}$ and state $\sf{\mathcal{S}}$, denoted ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$, occurred faithfully, without disclosing $\sf{\mathcal{S}}$ to $\sf{\mathcal{V}rf}$. Figure 1(b) highlights key differences between $\sf{PoX}$ and $\sf{PoSX}$ at a high level.
To realize $\sf{PoSX}$ on real-world IoT settings, we design and implement $\sf{SLAPP}$: a System-Level Approach for Poisoning Prevention. $\sf{SLAPP}$’s design and security rely on the commercially available TrustZone-M Security Extension, widely present even in low-end embedded devices: those based on ARM Cortex-M Micro-Controller Units (MCUs). This facilitates the immediate real-world implementation of $\sf{SLAPP}$ onto current IoT devices.
We show that $\sf{SLAPP}$ can support a wide range of data collection schemes, including FL and LDP, all with poisoning-free guarantees. Compared to prior data-centric mitigations in FL [50, 6, 33] or LDP [7], $\sf{SLAPP}$ offers two key benefits: it is agnostic to the underlying collection scheme (and implementation thereof), thereby capable of supporting both FL and LDP without changes in its trusted code base (TCB); it approaches the problem from a system security perspective, requiring no assumptions about the underlying data and its distribution. In summary, our anticipated contributions are the following:
  1. 1.
    New $\sf{PoSX}$ Security Notion: we define a new security primitive, called Proof of Stateful Execution ($\sf{PoSX}$). $\sf{PoSX}$ retains the same guarantees as classic $\sf{PoX}$ while addressing its limitations of inputless and stateless execution and maintaining privacy of underlying execution states vis-a-vis $\sf{\mathcal{V}rf}$.
  2. 2.
    Practical Poisoning Prevention: We develop $\sf{SLAPP}$: a design to realize $\sf{PoSX}$ that is applicable to resource-constrained embedded devices. We integrate $\sf{SLAPP}$ with FL and LDP implementations to support poison-free instantiations of these algorithms without loss of privacy.
  3. 3.
    Real-World Prototypes: To validate $\sf{SLAPP}$ and foster reproducibility, we provide three implementations, each utilizing distinct cryptographic schemes: symmetric, traditional asymmetric, and quantum-resistant asymmetric primitives. These implementations are prototyped and opensourced [3] using a real-world IoT development board: SAML11 Xplained Pro [36].
  4. 4.
    Evaluation: We conduct various experiments to assess $\sf{SLAPP}$’s efficiency. Our results demonstrate small runtime and memory overhead atop the baseline, in exchange for increased security and flexibility. Finally, we provide detailed case studies highlighting $\sf{SLAPP}$’s efficacy in thwarting poisoning attacks within FL/LDP.
II Background
II-A Privacy-Preserving Data Collection Schemes
In this section, we overview FL and LDP. We use the notation $\phi$ to indicate an empty/null variable.
Local Differential Privacy-based Data Collection (LDP-DC): The goal of this scheme is for an IoT sensor, $\sf{\mathcal{P}rv}$, to output a noisy sensor value ${\sf{\mathcal{O}}}$ such that it preserves $\epsilon$-LDP; informally speaking, $\epsilon$-LDP guarantees that ${\sf{\mathcal{O}}}$ leaks no information about its original value ${\sf{\mathcal{O}}}^{\prime}$ except with a small probability constrained by $\epsilon$. At the same time, collecting ${\sf{\mathcal{O}}}$ from a large number of $\sf{\mathcal{P}rv}$-s allows $\sf{\mathcal{V}rf}$ to obtain certain statistics of the collected sensor values with high confidence. Various LDP mechanisms have been proposed to achieve $\epsilon$-LDP in different settings. As an example case, we focus on the LDP mechanism called Basic RAPPOR [22] noting that the concepts introduced in this work can be generalized to other LDP-based mechanisms.
LDP-DC, as detailed in Algorithm 1, describes how $\sf{\mathcal{P}rv}$ performs data collection using Basic RAPPOR. First, $\sf{\mathcal{P}rv}$ collects a sensor reading ${\sf{\mathcal{O}}}^{\prime}$ by calling the sensor function $\sf{\mathcal{F}}$ on a given input $\sf{\mathcal{I}}$, i.e., ${\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$. Assuming ${\sf{\mathcal{O}}}^{\prime}$ can be represented using $k$-bit unsigned integer, $\sf{\mathcal{P}rv}$ next performs a unary encoding ($UE$) that transforms ${\sf{\mathcal{O}}}^{\prime}$ into a $2^{k}$-bit vector $b$ in which only the ${\sf{\mathcal{O}}}^{th}$ bit is set to 1 and other bits are 0.
Input : Sensor function $\sf{\mathcal{F}}$, Input to sensor function $\sf{\mathcal{I}}$, LDP parameters $f$, $p$ and $q$
Output : Noisy sensor output ${\sf{\mathcal{O}}}$
State : $PRR$ mapping $B$
1 Func ${\sf{\mathcal{F}}}$, ${\sf{\mathcal{I}}}$ $f$, $p$, $q$):
2       ${\sf{\mathcal{O}}}^{\prime}\leftarrow{\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$ $b\leftarrow UE({\sf{\mathcal{O}}}^{\prime})$ $b^{\prime}\leftarrow PRR(b,f,B)$ ${\sf{\mathcal{O}}}\leftarrow IRR(b^{\prime},p,q)$ return $\sf{\mathcal{O}}$
3
Algorithm 1 Implementation of LDP-DC on $\sf{\mathcal{P}rv}$
Next, on input $b$ and LDP parameter $f$, $\sf{\mathcal{P}rv}$ invokes Permanent Randomized Response ($PRR$) function and produces a $2^{k}$-bit noisy vector $b^{\prime}$ as output, where $b^{\prime}_{i}$ – the $i^{th}$ bit of $b^{\prime}$ – is computed as:
$\footnotesize b^{\prime}_{i}=\begin{cases}1,&\text{with probability }\frac{f}{2}\\ 0,&\text{with probability }\frac{f}{2}\\ b_{i},&\text{with probability }1-f\\ \end{cases}$ (1)
Once $b^{\prime}$ is generated, $\sf{\mathcal{P}rv}$ caches an input-output mapping ($b$, $b^{\prime}$) to a local state variable $B$ (i.e., $B[b]=b^{\prime}$) and, for all future encounters of the same input $b$, returns $B[b]$ without recomputing the entire $PRR$ function.
Finally, on input $b^{\prime}$ and LDP parameters $p$ and $q$, $\sf{\mathcal{P}rv}$ executes Instantaneous Randomized Response ($IRR$) function that returns a $2^{k}$-bit binary vector ${\sf{\mathcal{O}}}$ to $\sf{\mathcal{V}rf}$ such that:
$\footnotesize\mathbb{P}({\sf{\mathcal{O}}}_{i}=1)=\begin{cases}p,&\text{if }b^{\prime}_{i}=1\\ q,&\text{otherwise}\end{cases}$ (2)
Using LDP-DC, $\sf{\mathcal{V}rf}$ can estimate $\tilde{f}_{x}$, the frequency of sensor value $x$ by:
$\footnotesize\tilde{f}_{x}=\dfrac{c_{x}-(q+\frac{1}{2}fp-\frac{1}{2}fq)n}{(1-f)(p-q)n}$ (3)
where $c_{x}$ represents the number of reports that have $x^{th}$ bit set and $n$ is the number of reports received by $\sf{\mathcal{V}rf}$. We refer the interested reader to the Basic RAPPOR paper [22] for details on how to select $f,p,q$ to satisfy $\epsilon$-LDP.
Importantly, privacy in this scheme requires $\sf{\mathcal{V}rf}$ to be oblivious to $B$; otherwise, $\sf{\mathcal{V}rf}$ can reverse the $PRR$ operation to recover the original sensor reading ${\sf{\mathcal{O}}}^{\prime}$. Also, this scheme requires multiple input arguments as part of its execution (i.e., LDP parameters $f$, $p$, and $q$). As noted earlier, however, standard PoX does not support execution using external input arguments or $\sf{\mathcal{P}rv}$ local states that must be oblivious to $\sf{\mathcal{V}rf}$
Input : Sesnor function $\sf{\mathcal{F}}$, Input to sensor function $\sf{\mathcal{I}}$
Output : $\phi$
State : Local dataset $D$
1 Func $\sf{\mathcal{F}}$, $\sf{\mathcal{I}}$):
2       ${\sf{\mathcal{O}}}\leftarrow{\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$ $D$.append($\sf{\mathcal{O}}$)
3
Input : Globally trained weights $W$, number of epochs $t$, learning rate $\alpha$
Output : Locally trained weights ${\sf{\mathcal{O}}}$
State : $D$
4 Func Train($W$, $t$, $\alpha$):
5       for $k\leftarrow 1$ to $t$ do
             // $\nabla$ is a gradient function
6             $W\leftarrow W-\alpha\cdot\nabla(W;D)$
7       end for
8      ${\sf{\mathcal{O}}}\leftarrow W$ return $\sf{\mathcal{O}}$
9
Algorithm 2 Implementation of FL-DC on $\sf{\mathcal{P}rv}$
Federated Learning-based data collection (FL-DC): Contrary to LDP-DC, FL-DC [30] requires $\sf{\mathcal{P}rv}$ to send a locally trained machine learning (ML) partial model to $\sf{\mathcal{V}rf}$ instead of sensor readings. This keeps raw sensor data local to $\sf{\mathcal{P}rv}$ and not directly accessible by $\sf{\mathcal{V}rf}$.
A typical FL-DC consists of two phases depicted in Algorithm 2. In the first phase, $\sf{\mathcal{P}rv}$ invokes the Sense-Store function to collect a raw sensor reading and store it in a local list $D$. This phase can be repeated multiple times over a certain period to collect more training data on $\sf{\mathcal{P}rv}$. Once a sufficient amount of training data is gathered, $\sf{\mathcal{P}rv}$ receives globally trained weights $W$ from $\sf{\mathcal{V}rf}$ along with the training parameters ($t$ and $\alpha$). It then triggers the Train function that utilizes $W$ as a base model to train the data in $D$. This function outputs the locally trained weights ${\sf{\mathcal{O}}}$ without disclosing $D$ to $\sf{\mathcal{V}rf}$.
After collecting weights ${\sf{\mathcal{O}}}$-s from multiple devices, $\sf{\mathcal{V}rf}$ can aggregate them using several methods, e.g., FedAvg [30] averages the received weights and sets the result as the new global weights: $W\leftarrow W+\eta\cdot\sum_{i=1}^{m}({\sf{\mathcal{O}}}_{i}-W)/m$ for some global learning rate $\eta$.
Similar to LDP-DC, executing FL-DC functions relies on both input arguments (i.e., training input arguments $W$, $t$ and $\alpha$, as well as $\sf{\mathcal{P}rv}$-local state $D$). Consequently, FL-DC integration with classic $\sf{PoX}$ faces the same challenges as LDP-DC.
II-B ARM TrustZone-M Security Extensions
ARM TrustZone-M is a hardware security extension that enables a trusted execution environment (TEE) in ARM Cortex-M MCUs commonly used in low-cost and energy-efficient IoT applications. TrustZone divides software states into two isolated worlds: Secure and Non-Secure. The Non-Secure world contains and executes (untrusted) application software while security-critical (trusted) software is stored and runs in the Secure world.
In particular, we leverage two security properties of ARM TrustZone-M in this work:
  • Hardware-enforces World Isolation. TrustZone-M ensures complete isolation of these worlds by implementing several hardware controls (i.e., TZASC, TZMA modules) to enforce access control to hardware resources (e.g., program and data memory, peripherals) for these two worlds. With isolation in place, TrustZone ensures the Non-Secure World is unable to access any code and data located in the Secure World. This assures that the Secure world remains secure even if an adversary can fully modify or compromise the Non-Secure world software state.
  • Controlled Invocation. In TrustZone-M, the only legal way for the non-secure world to access a function inside the secure world is by issuing the Security Monitor Call (SMC) instruction. This instruction provides a secure context switch to the secure world. In addition, the SMC instruction ensures correct invocation of the secure world’s functions, i.e., it always invokes this function at its first legal instruction while prohibiting a jump to the middle of the function.
III System Model and Assumptions
III-A Network and Usage Model
We consider an IoT setting as shown earlier in Figure 1, consisting of two entity types: one $\sf{\mathcal{V}rf}$ and multiple $\sf{\mathcal{P}rv}$-s. $\sf{\mathcal{P}rv}$ is a resource-constrained sensor deployed in a physical space of interest, e.g., a smart home, office, or factory. $\sf{\mathcal{V}rf}$ is a remote service provider that orchestrates these $\sf{\mathcal{P}rv}$-s. As an example, $\sf{\mathcal{P}rv}$-s could be smart light bulbs that are used in many smart homes and can be controlled by the end-user through Philips $\sf{\mathcal{V}rf}$’s application services; or Samsung could act as $\sf{\mathcal{V}rf}$ that provides a service for the end-user to command all SmartThings-compatible devices.
Besides offering this service, $\sf{\mathcal{V}rf}$ wishes to collect sensor data generated by $\sf{\mathcal{P}rv}$-s to further improve the service performance or enhance the user experience. In the context of this work, $\sf{\mathcal{V}rf}$ has the option to employ LDP-DC or FL-DC as its preferred data collection scheme, depending on the desired outcome and privacy considerations. As poisoning attacks could sabotage the collection outcome, $\sf{\mathcal{V}rf}$ also aims to detect and prevent such attacks to safeguard the authenticity of the outcomes.
III-B Adversary Model
We consider an $\sf{\mathcal{A}dv}$ who can modify/compromise $\sf{\mathcal{P}rv}$’s application software at will. Once compromised, $\sf{\mathcal{A}dv}$ can access, modify, or erase any code or data in $\sf{\mathcal{P}rv}$ unless explicitly protected by hardware-enforced access control rules. Consequently, $\sf{\mathcal{A}dv}$ may use this ability to perform poisoning attacks by corrupting a sensor function $\sf{\mathcal{F}}$ or its execution to spoof arbitrary results sent to $\sf{\mathcal{V}rf}$. Invasive hardware-based/physical attacks (e.g., fault-injection attacks or physical hardware manipulation) are out of scope in this work, as they require orthogonal tamper-proofing techniques [47].
We also assume $\sf{\mathcal{A}dv}$ has full control over the communication channel between $\sf{\mathcal{P}rv}$ and $\sf{\mathcal{V}rf}$. They may perform any network-based attacks, e.g., reading, modifying, replaying, or dropping any message sent from/to $\sf{\mathcal{P}rv}$.
Finally, in line with the standard LDP and FL $\sf{\mathcal{A}dv}$ models, we also consider the possibility of malicious $\sf{\mathcal{V}rf}$. In the latter, $\sf{\mathcal{A}dv}$’s goal is to learn sensitive data on $\sf{\mathcal{P}rv}$, which could be accomplished by exploiting weaknesses in the underlying data collection scheme.
III-C Device Model
$\sf{\mathcal{P}rv}$-s are small embedded/IoT devices equipped with TrustZone-M, e.g., ARM embedded devices running on Cortex-M23/33 MCUs, which are optimized for low-cost and energy efficiency. Following the PoX assumption [13], we assume that the function whose execution is being proven ($\sf{\mathcal{F}}$, i.e., the code implementing the data collection task according to the underlying scheme) is correct and contains no implementation bug that can lead to run-time exploits within itself. In practice, $\sf{\mathcal{V}rf}$ can employ various pre-deployment vulnerability detection techniques to fulfill this requirement [10].
We also adhere to standard TEE-based security assumptions, i.e., we assume the small TCB implementing the $\sf{PoSX}$ RoT located inside TrustZone’s Secure World is trusted and TrustZone hardware modules are implemented correctly such that $\sf{\mathcal{A}dv}$ cannot modify this RoT implementation or violate any security guarantees implemented by the Secure World-resident code. The latter implies the existence of secure persistent storage, exclusively accessible by the Secure World and unmodifiable when the device is offline. This storage is used to store our TCB along with a counter-based challenge $c$ and two cryptographic keys: $\sf{sk_{{\sf{\mathcal{P}rv}}}}$ and $\sf{pk_{{\sf{\mathcal{V}rf}}}}$, where $\sf{sk_{{\sf{\mathcal{P}rv}}}}$ corresponds to $\sf{\mathcal{P}rv}$ private key, whose public counterpart $\sf{pk_{{\sf{\mathcal{P}rv}}}}$ is known to $\sf{\mathcal{V}rf}$. Similarly, $\sf{pk_{{\sf{\mathcal{V}rf}}}}$ denotes $\sf{\mathcal{V}rf}$ public key with its private counterpart $\sf{sk_{{\sf{\mathcal{P}rv}}}}$ securely managed by $\sf{\mathcal{V}rf}$. In TrustZone-M implementations, secure persistent memory is supported by a standard secure boot architecture [4] with the physical memory storing cryptographic keys being physically inaccessible through I/O interfaces (e.g., USB/J-TAG, etc.).
Finally, we assume these keys are correctly distributed to $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$ out-of-band, e.g., by physically fusing these keys on $\sf{\mathcal{P}rv}$ during manufacture time or using any key-provisioning mechanism [37] after $\sf{\mathcal{P}rv}$ deployment.
IV $\sf{PoSX}$ and Associated Definitions
To define $\sf{PoSX}$ security goal, we start by revisiting classic $\sf{PoX}$ guarantees and their limitations. Then we formulate auxiliary notions that address each limitation and in conjunction imply $\sf{PoSX}$ end-to-end goal.
{definition}
[PoX Security]Let $\sf{\mathcal{F}}$ represent an arbitrary software function (code) execution of which is requested by $\sf{\mathcal{V}rf}$ on $\sf{\mathcal{P}rv}$, producing output ${\sf{\mathcal{O}}}$. A protocol is considered PoX-secure if and only if the protocol outputting $\top$ implies:
  1. (i)
    ${\sf{\mathcal{F}}}$ code (as defined by $\sf{\mathcal{V}rf}$) executes atomically and completely between ${\sf{\mathcal{V}rf}}$ sending a request and receiving the response, and
  2. (ii)
    $\sf{\mathcal{O}}$ is a direct outcome of this execution of ${\sf{\mathcal{F}}}()$.
Definition IV states the classic $\sf{PoX}$ security notion, as described in [13]. A violation of conditions (i) or (ii) in Definition IV must be detectable by $\sf{\mathcal{V}rf}$, resulting in a protocol abort (i.e., by outputting $\bot$). As discussed in [13], this notion can be used to construct “sensors that cannot lie” irrespective of compromised software states. However, it only supports self-contained IoT applications that are independent of $\sf{\mathcal{V}rf}$-defined inputs or pre-computed states due to two limitations:
L1 – Lack of input validation. Definition IV considers inputless $\sf{\mathcal{F}}$ functions. This is because classic $\sf{PoX}$ does not support verification of the origin and integrity of inputs received by $\sf{\mathcal{P}rv}$. This limitation is significant for applications where $\sf{\mathcal{V}rf}$ must provide input $\sf{\mathcal{I}}$ as part of $\sf{\mathcal{F}}$ execution on $\sf{\mathcal{P}rv}$ (e.g., FL/LDP cases discussed in Sec. II-A).
L2 – No state preservation across $\sf{PoX}$ instances. Definition IV only supports $\sf{PoX}$ of stateless $\sf{\mathcal{F}}$ functions. In other words, $\sf{\mathcal{F}}$ must rely solely on data generated/acquired within its current execution instance and must not depend on $\sf{\mathcal{P}rv}$ states (denoted $\sf{\mathcal{S}}$) produced by prior executions. Similar to L1 case, a $\sf{PoX}$ protocol satisfying Definition IV provides no guarantee or validation to the correct use of some pre-existent/expected state $\sf{\mathcal{S}}$ in $\sf{\mathcal{P}rv}$. Thus, attacks that tamper with $\sf{\mathcal{S}}$ in between subsequent $\sf{PoX}$ instances may result in illegal alteration of the end result $\sf{\mathcal{O}}$.
Remark: note that when $\sf{\mathcal{S}}$ contains only public information, L2 can be obviated by L1 by making $\sf{\mathcal{S}}$ a part of the authenticated output of a $\sf{PoX}$ instance and used as a $\sf{\mathcal{V}rf}$-defined input to $\sf{\mathcal{F}}$ in a subsequent $\sf{PoX}$ instance. However, when $\sf{\mathcal{S}}$ contains information that must remain hidden from $\sf{\mathcal{V}rf}$ (the case of our target applications – recall Sec. II-A), consistency of $\sf{\mathcal{S}}$ must be ensured locally at $\sf{\mathcal{P}rv}$, making L1 and L2 independent challenges.
To address these limitations systematically, we introduce two new $\sf{PoX}$-related security notions. As these notions may be of independent interest, we first present them separately and finally compose them into an end-to-end $\sf{PoSX}$ goal.
{definition}
[IV-$\sf{PoX}$ Security]Let $\sf{\mathcal{F}}$ represent an arbitrary software function (code) and ${\sf{\mathcal{I}}}$ represent an input, both defined by $\sf{\mathcal{V}rf}$. Let ${\sf{\mathcal{O}}}$ represent the output and $\sigma$ denote a proof of ${\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$ execution produced by $\sf{\mathcal{P}rv}$. A protocol is considered IV-$\sf{PoX}$-secure if and only if the protocol outputting $\top$ implies:
  1. (i)
    ${\sf{\mathcal{F}}}$ executes with input arguments ${\sf{\mathcal{I}}}$, atomically and completely between ${\sf{\mathcal{V}rf}}$ sending the request and receiving $\sigma$, and
  2. (ii)
    $\sf{\mathcal{O}}$ is a direct outcome of this ${\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$ execution
To overcome L1, we present the notion of input-validating $\sf{PoX}$ (or IV-$\sf{PoX}$), as shown in Definition IV. In IV-$\sf{PoX}$, $\sf{\mathcal{V}rf}$ aims to execute $\sf{\mathcal{F}}$ with its own provided input $\sf{\mathcal{I}}$. Thus, the proof generated by $\sf{\mathcal{P}rv}$, $\sigma$, must not only validate atomic and complete execution of $\sf{\mathcal{F}}$ but also that it was invoked with the correct input requested by $\sf{\mathcal{V}rf}$ (as captured in condition (i) of Definition IV). Thus, an IV-$\sf{PoX}$ protocol must ensure $\sf{\mathcal{O}}$ authenticity concerning $\sf{\mathcal{I}}$, i.e., $\sf{\mathcal{O}}$ is generated by executing $\sf{\mathcal{F}}$ correctly using the expected input $\sf{\mathcal{I}}$.
{definition}
[SP-$\sf{PoX}$ Security]Let $\sf{\mathcal{F}}$ represent an arbitrary software function (code) requested by $\sf{\mathcal{V}rf}$ to run on $\sf{\mathcal{P}rv}$ with state ${\sf{\mathcal{S}}}$, where $\sf{\mathcal{S}}$ was produced by some prior $\sf{PoX}$ on $\sf{\mathcal{P}rv}$ but is oblivious to $\sf{\mathcal{V}rf}$. Let ${\sf{\mathcal{O}}}$ represent the output and $\sigma$ denote a proof of ${\sf{\mathcal{F}}}$ execution produced by $\sf{\mathcal{P}rv}$. A protocol is considered SP-$\sf{PoX}$-secure if and only if the protocol outputting $\top$ implies:
  1. (i)
    ${\sf{\mathcal{F}}}$ executes atomically and completely between ${\sf{\mathcal{V}rf}}$ sending the request and receiving $\sigma$ with $\sf{\mathcal{P}rv}$ state corresponding to $\sf{\mathcal{S}}$ when $\sf{\mathcal{F}}$ execution starts (denote this execution by ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}()$), and
  2. (ii)
    $\sf{\mathcal{O}}$ is a direct outcome of this ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}()$ execution, and
  3. (iii)
    $\sf{\mathcal{V}rf}$ cannot infer the value of $\sf{\mathcal{S}}$ beyond what is revealed by $\sf{\mathcal{O}}$, and
  4. (iv)
    $\sf{\mathcal{S}}$ was not modified between the current ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}()$ execution and the prior $\sf{\mathcal{V}rf}$-authorized $\sf{PoX}$.
To address L2, we specify the State-Preserving $\sf{PoX}$ (SP-$\sf{PoX}$) notion in Definition IV. Similar to IV-$\sf{PoX}$ security, the SP-$\sf{PoX}$ notion specifies the first two conditions to ensure that, in addition to atomic $\sf{\mathcal{F}}$ execution, $\sigma$ also conveys two critical aspects: (1) correct use of $\sf{\mathcal{S}}$ during $\sf{\mathcal{F}}$ execution and (2) dependence of $\sf{\mathcal{O}}$ on $\sf{\mathcal{F}}$ and $\sf{\mathcal{S}}$. In addition, it requires $\sf{\mathcal{S}}$ privacy vis-a-vis $\sf{\mathcal{V}rf}$ and prohibits $\sf{\mathcal{S}}$ modification in between subsequent $\sf{PoX}$ instances.
{definition}
[$\sf{PoSX}$ Security] A scheme is $\sf{PoSX}$-secure if and only if it satisfies both IV-$\sf{PoX}$ (Def. IV) and SP-$\sf{PoX}$ (Def. IV) Security.
Finally, Definition IV combines IV-$\sf{PoX}$ and SP-$\sf{PoX}$ to state the goal of $\sf{PoSX}$-Security.
V $\sf{SLAPP}$: Realizing $\sf{PoSX}$ with TrustZone-M
V-A Overview of $\sf{SLAPP}$ Workflow
Building on TrustZone-M hardware-enforced world isolation (recall Sec. II-B), our approach is to implement $\sf{SLAPP}$’s RoT in and execute it from the Secure World. Meanwhile, normal applications are untrusted (hereby referred to as untrusted software) and reside in the Non-Secure World.
$\sf{SLAPP}$ implements three Secure World functions: Execute, CheckState, and SetState. Execute serves as the main call to execute ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ on $\sf{\mathcal{P}rv}$ and compute a proof of this stateful execution. CheckState and SetState are used to authenticate state $\sf{\mathcal{S}}$ used in Execute.
$\sf{SLAPP}$ workflow is depicted in Figure 3. At a high level, it enforces the following operation sequence upon receiving a $\sf{PoSX}$ request from $\sf{\mathcal{V}rf}$.:
  • Upon being called by the Non-Secure World, Execute authenticates $\sf{\mathcal{V}rf}$ request in ❹.
  • Following successful authentication, it starts executing ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ atomically in Non-Secure World, ❺.
  • In order to access $\sf{\mathcal{S}}$ in ❻, ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ must call CheckState function in the Secure World. CheckState takes one input argument, representing the current $\sf{\mathcal{S}}$ value. Its task is to authenticate this value by matching it against the latest benign $\sf{\mathcal{S}}$ value from a prior execution stored within the Secure World.
  • Similarly, after $\sf{\mathcal{S}}$ is modified in ❻, execution is trapped into the Secure World via the SetState function. SetState accepts one input argument: the new $\sf{\mathcal{S}}$ value. It is responsible for committing and securely maintaining the latest benign $\sf{\mathcal{S}}$ value in the Secure World.
  • Once ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ execution completes, yielding output $\sf{\mathcal{O}}$, the control returns to Execute in ❼. Then, Execute computes $\sigma$ indicating an authenticated proof of this execution in ❽ and returns ($\sf{\mathcal{O}}$, $\sigma$) to the Non-Secure World in ❾, which in turn forwards them to $\sf{\mathcal{V}rf}$.
With ($\sf{\mathcal{O}}$, $\sigma$), $\sf{\mathcal{V}rf}$ can determine if ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ was executed in $\sf{\mathcal{P}rv}$’s Non-Secure World, if $\sf{\mathcal{O}}$ is an authentic result of this execution, and if $\sf{\mathcal{S}}$ was preserved since the prior execution. Importantly, $\sf{SLAPP}$ assures that $\sigma$ is not computable unless the aforementioned operation sequence is observed.
Figure 3: Overview of $\sf{SLAPP}$ workflow
V-B $\sf{SLAPP}$ in Detail
Protocol V-B details the $\sf{\mathcal{V}rf}$leftrightarrow $\sf{\mathcal{P}rv}$ interation in $\sf{SLAPP}$
Phase 1: Request Generation
An instance of $\sf{SLAPP}$ protocol starts when $\sf{\mathcal{V}rf}$ generates a $\sf{PoSX}$ request, in Step 1, comprising: (i) an identifier ${\sf{\mathcal{F}}}$ for which function to execute; (ii) input arguments $\sf{\mathcal{I}}$; and (iii) a monotonically-increasing counter-based challenge $c_{\sf{\mathcal{V}rf}}$. This request instructs $\sf{\mathcal{P}rv}$ to execute ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ in the Non-Secure World on $\sf{\mathcal{P}rv}$ and return the result to $\sf{\mathcal{V}rf}$ along with proof of this expected execution. Following this, $\sf{\mathcal{V}rf}$ signs this request (via Sign function), producing token $\sigma_{\sf{\mathcal{V}rf}}$, and attaches this token to the request before sending it to $\sf{\mathcal{P}rv}$ in Step 2 of Protocol V-B.
Phase 2: Preparation for ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ Execution
In Step 3, untrusted software in $\sf{\mathcal{P}rv}$ must invoke Execute passing the $\sf{PoSX}$ request as a parameter. In step 4, Execute (i) checks whether the request counter $c_{\sf{\mathcal{V}rf}}$ is larger than a local counter $c$ maintained by the Secure World to ensure freshness of the $\sf{PoSX}$ request and prevent replayed attacks; (ii) verifies $\sigma_{\sf{\mathcal{V}rf}}$ to confirm authenticity of the request; and (iii) checks that no other $\sf{SLAPP}$ instance is active by examining $exec$ flag maintained in the secure world. If any of the checks fail, the process is aborted which implies the inability to produce the end proof $\sigma$. The same applies if Execute is never called.
If all checks succeed, Execute updates the local counter with $c$ with $c_{\sf{\mathcal{V}rf}}$ and initializes three Secure World flags in Step 5:
  • $exec$ is set to $\top$, indicating an active $\sf{SLAPP}$ instance.
  • $stateChecked$ is set to $\bot$, indicating the status of $\sf{\mathcal{S}}$ authenticity during $\sf{\mathcal{F}}$ execution.
  • $stateUsed$ is set to $\bot$, indicating the status of $\sf{\mathcal{S}}$ access during $\sf{\mathcal{F}}$ execution.
In Step 6, Execute prepares $\sf{\mathcal{P}rv}$ for the upcoming execution of $\sf{\mathcal{F}}$ by disabling all interrupts and taking a snapshot ($h$) as a hash digest reflecting the states of the Non-Secure World’s binary code in program memory ($PMEM$) and the parameters received in the $\sf{PoSX}$ request. It then calls ${\sf{\mathcal{F}}}$ on input $\sf{\mathcal{I}}$ in the Non-Secure World (Step 7).
To leverage $\sf{SLAPP}$, the implementation of $\sf{\mathcal{F}}$ must:
  • B1:
    At the start of its execution, validate relevant state $\sf{\mathcal{S}}$ authenticity by calling CheckState.
  • B2:
    At the end of its execution, commit the latest $\sf{\mathcal{S}}$ value to the Secure World by invoking SetState.
  • B3:
    Throughout its execution, $\sf{\mathcal{F}}$ must never enable interrupts.
Phase 3: ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ Execution
From B1, it follows that a call to $\sf{\mathcal{F}}$, in Step 8, triggers CheckState of current $\sf{\mathcal{S}}$. In Step 9, CheckState, executed in the Secure World, sets $stateUsed$ to $\top$ and proceeds to verify authenticity of the current $\sf{\mathcal{S}}$ value, $s$, by comparing $s$ with the latest benign $\sf{\mathcal{S}}$ value, ${\sf{\mathcal{S}}}_{sec}$, stored in the Secure World. Only if they match, CheckState ascertains $s$ authenticity setting $stateChecked$ to $\top$. Execution of $\sf{\mathcal{F}}$ is then resumed. From B2, SetState is invoked at the end of $\sf{\mathcal{F}}$ execution to update ${\sf{\mathcal{S}}}_{sec}$ with the new state $s$ if and only if ($stateChecked=exec=\top$). To optimize storage, especially when $|s|$ is large, SetState can update ${\sf{\mathcal{S}}}_{sec}$ with a hash of $s$, as shown in Step 11. This storage optimization comes with the expense of additional runtime overhead for hash computations in CheckState and SetState. We discuss this time-space trade-off further in Section VII.
$\sf{\mathcal{F}}$ execution completes producing the output $\sf{\mathcal{O}}$ and handing the control to Execute with $\sf{\mathcal{O}}$ as an input in Step 12.
Phase 4: Proof Generation
Execute examines $exec$, $stateUsed$, and $stateChecked$ flags to determine the occurrence of the CheckState $\rightarrow$ SetState sequence. If this sequence is maintained during $\sf{\mathcal{F}}$ execution, Execute proceeds to compute the proof $\sigma$ in Step 13 by signing $h$ and $\sf{\mathcal{O}}$ using $\sf{\mathcal{P}rv}$’s private key ${\sf{sk_{{\sf{\mathcal{P}rv}}}}}$. Execute resets all Secure World flags before returning to the Non-Secure World with $\sf{\mathcal{O}}$ and $\sigma$, in Step 15. They are then transmitted to $\sf{\mathcal{V}rf}$, in Step 16.
Protocol 1: $\sf{SLAPP}$ Protocol
Verifier ($\sf{\mathcal{V}rf}$)
  1. (1)
    Generate an authenticated $\sf{PoSX}$ request:
    $\sigma_{\sf{\mathcal{V}rf}}\leftarrow\textsf{Sign}({\sf{sk_{{\sf{\mathcal{V}rf}}}}},\textsf{H}({\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}}))$
  2. (2)
    Send ($\sigma_{\sf{\mathcal{V}rf}}$, $\sf{\mathcal{F}}$, $\sf{\mathcal{I}}$, $c_{\sf{\mathcal{V}rf}}$) to $\sf{\mathcal{P}rv}$
Prover ($\sf{\mathcal{P}rv}$)
In Non-Secure World:
  1. (3)
    Call Execute in Secure World with the received request
In Execute function, Secure World:
  1. (4)
    Authenticate the request and abort if $r=\bot$ or $exec=\top$:
    $r\leftarrow(c_{\sf{\mathcal{V}rf}}>c)\land\textsf{Verify}({\sf{pk_{{\sf{\mathcal{V}rf}}}}},\sigma_{\sf{\mathcal{V}rf}},{\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}})$
  2. (5)
    Update counter: $c\leftarrow c_{\sf{\mathcal{V}rf}}$ and initialize: $exec\leftarrow\top$, $stateUsed\leftarrow\bot$ $stateChecked\leftarrow\bot$
  3. (6)
    Disable interrupts and measure $PMEM$ and $\sf{\mathcal{V}rf}$ request:
    $h\leftarrow\textsf{H}(PMEM,{\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}})$
  4. (7)
    Call ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ in Non-Secure World
In $\sf{\mathcal{F}}$ function, Non-Secure World:
  1. (8)
    Run ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$. Pass control to the Secure World when CheckState($s$) is called
In CheckState function, Secure World:
  1. (9)
    Perform an integrity check on $s$, store the result to $p$ and return to Non-Secure World: $stateChecked\leftarrow(\textsf{H}(s)\stackrel{{\scriptstyle?}}{{=}}{\sf{\mathcal{S}}}_{sec})$. Also, set $stateUsed\leftarrow\top$.
In $\sf{\mathcal{F}}$ function, Non-Secure World:
  1. (10)
    Continue with ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ execution. Pass control to the Secure World when SetState($s$) is called
In SetState function, Secure World:
  1. (11)
    Securely set ${\sf{\mathcal{S}}}_{sec}$ based on input $s$ and return to Non-Secure World:
    ${\sf{\mathcal{S}}}_{sec}\leftarrow\textsf{H}(s)\text{ if }(stateChecked\land exec)$
In $\sf{\mathcal{F}}$ function, Non-Secure World:
  1. (12)
    Continue with ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ until the execution is completed, producing output $\sf{\mathcal{O}}$, and then return to its caller, Execute, with $\sf{\mathcal{O}}$
In Execute function, Secure World:
  1. (13)
    Abort if $(exec~{}\land stateUsed~{}\land\neg stateChecked)$. Otherwise, include $\sf{\mathcal{O}}$ to the measurement and compute the proof:
    $\sigma\leftarrow\textsf{Sign}({\sf{sk_{{\sf{\mathcal{P}rv}}}}},\textsf{H}(h,{\sf{\mathcal{O}}}))$
  2. (14)
    Reset all flags: $exec\leftarrow\bot,stateChecked\leftarrow\bot,stateUsed\leftarrow\bot$
  3. (15)
    Enable interrupts and return with $({\sf{\mathcal{O}}},\sigma)$
In Non-Secure World:
  1. (16)
    Forward $({\sf{\mathcal{O}}},\sigma)$ to $\sf{\mathcal{V}rf}$
Verifier ($\sf{\mathcal{V}rf}$)
  1. (17)
    Increment $c_{\sf{\mathcal{V}rf}}$ and validate $\sigma$ by:
    $r\leftarrow\textsf{ValidatePoSX}({\sf{pk_{{\sf{\mathcal{P}rv}}}}},\sigma,PMEM^{\prime},{\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}},{\sf{\mathcal{O}}})$
    The protocol outputs $r$.
Phase 5: Proof Validation
Upon receiving $\sf{\mathcal{O}}$ and $\sigma$, $\sf{\mathcal{V}rf}$ increments $c_{\sf{\mathcal{V}rf}}$ and performs the $\sf{PoSX}$ verification in Step 17 by:
  • Checking validity of $\sigma$. As $\sf{\mathcal{V}rf}$ possesses the expected binary of $\sf{\mathcal{P}rv}$’s Non-Secure World, $PMEM^{\prime}$, this verification involves checking $\sigma$ against $PMEM^{\prime}$, $\sf{\mathcal{F}}$, $\sf{\mathcal{I}}$, $c_{\sf{\mathcal{V}rf}}$ and $\sf{\mathcal{O}}$ using $\sf{\mathcal{P}rv}$’s public key ${\sf{pk_{{\sf{\mathcal{P}rv}}}}}$, i.e.:
    $\textsf{Verify}({\sf{pk_{{\sf{\mathcal{P}rv}}}}},\sigma,PMEM^{\prime},{\sf{\mathcal{F}}},{\sf{\mathcal{I}}},{\sf{\mathcal{O}}},c_{\sf{\mathcal{V}rf}})\stackrel{{\scriptstyle?}}{{=}}\top$
  • Inspect ${\sf{\mathcal{F}}}$ binary to ensure that it adheres to the expected behaviors: B1, B2 and B3.
Finally, $\sf{SLAPP}$ protocol is considered successful and thus outputs $\top$ if it passes both checks; it aborts with $\bot$ otherwise.
VI $\sf{PoSX}$ Security Analysis
Our security argument is based on the following properties:
P1 - Request Verification. In $\sf{SLAPP}$, $\sf{\mathcal{P}rv}$’s TCB always verifies freshness and authenticity of a $\sf{PoSX}$ request before executing $\sf{\mathcal{F}}$ and generating the proof $\sigma$. This prevents $\sf{\mathcal{A}dv}$ from exploiting forged or replayed requests to manipulate the protocol outcome. See step 4 of Protocol V-B.
P2 - Input Validation. A valid $\sigma$ serves as authentication for the correct usage of $\sf{\mathcal{I}}$ during $\sf{\mathcal{F}}$ execution. This prevents $\sf{\mathcal{A}dv}$ from feeding malicious input to $\sf{\mathcal{F}}$ while still succeeding in the $\sf{SLAPP}$ protocol. Since $\sf{\mathcal{I}}$ is included in a $\sf{PoSX}$ request, this is implied by P1 and the fact that $\sf{\mathcal{I}}$ is directly used by $\sf{SLAPP}$ to invoke ${\sf{\mathcal{F}}}$ in step 7 of Protocol V-B.
P3 - State Privacy. $\sf{SLAPP}$ protects privacy of $\sf{\mathcal{S}}$ from $\sf{\mathcal{V}rf}$ since the only information $\sf{\mathcal{V}rf}$ receives are $\sf{\mathcal{O}}$ and $\sigma$. In cases that $\sf{\mathcal{S}}$ can be leaked through $\sf{\mathcal{O}}$, $\sf{\mathcal{P}rv}$ has the option to refuse to send it over to $\sf{\mathcal{V}rf}$ by simply terminating the protocol. Moreover, since $\sigma$ is not a function of $\sf{\mathcal{S}}$, it leaks no information about $\sf{\mathcal{S}}$ to $\sf{\mathcal{V}rf}$.
P4 - State Authenticity. In $\sf{SLAPP}$, the successful completion of a protocol instance guarantees the $\sf{\mathcal{S}}$ value, stored in the Non-Secure World, is authentic, i.e., it can only be modified by $\sf{\mathcal{V}rf}$-approved software during a protocol instance and remains unchanged between consecutive instances. This assurance comes from two observations:
First, ${\sf{\mathcal{S}}}_{sec}$ always corresponds to the latest benign $\sf{\mathcal{S}}$ value because: (1) ${\sf{\mathcal{S}}}_{sec}$ cannot be updated outside a protocol instance due to the check of $exec$ flag in Step 11; and (2) ${\sf{\mathcal{S}}}_{sec}$ cannot be influenced by forged or replayed $\sf{PoSX}$ requests since the request is always authenticated in Step 4 before $exec$ can be set in Step 5.
Second, any unauthorized modification to $\sf{\mathcal{S}}$ outside a protocol instance is detected in the subsequent instance by CheckState due to a mismatch between the $\sf{\mathcal{S}}$ value and ${\sf{\mathcal{S}}}_{sec}$. As a successful $\sf{SLAPP}$ protocol implies a successful check from CheckState, $\sf{\mathcal{S}}$ must be equal to ${\sf{\mathcal{S}}}_{sec}$, and, according to the first observation, must contain the latest authentic $\sf{\mathcal{S}}$ value. Moreover, $\sf{\mathcal{A}dv}$ may attempt to tamper with $\sf{\mathcal{S}}$ while $\sf{\mathcal{F}}$ is executing. However, doing so requires modification to $PMEM$, which would result in a mismatch with $PMEM^{\prime}$ during the proof validation in Step 17.
P5 - Atomic Execution. $\sf{\mathcal{F}}$ execution must occur atomically; otherwise, the protocol must fail. This is required to prevent $\sf{\mathcal{A}dv}$ from interrupting $\sf{\mathcal{F}}$ execution to tamper with its data and execution flows, influencing the outcome. Step 6 realizes this requirement by disabling all interrupts before ${\sf{\mathcal{F}}}$ invocation. We explain how this requirement can be relaxed in Section VII.
P6 - Output Authenticity. A successful $\sf{SLAPP}$ protocol indicates to $\sf{\mathcal{V}rf}$ that $\sf{\mathcal{O}}$ is authentic and generated by executing $\sf{\mathcal{F}}$ atomically with the $\sf{\mathcal{V}rf}$-specified $\sf{\mathcal{I}}$ and the correct state $\sf{\mathcal{S}}$. $\sf{SLAPP}$ satisfies this property since P1 guarantees that $\sf{\mathcal{F}}$ is always invoked with authentic $\sf{\mathcal{I}}$ and P5 enforces $\sf{\mathcal{F}}$ to execute without interruptions before immediately returning to Secure World with $\sf{\mathcal{O}}$. This leaves no opportunities for untrusted software to interrupt $\sf{\mathcal{F}}$ execution to tamper with $\sf{\mathcal{O}}$. $\sf{\mathcal{A}dv}$ attempt to change $\sf{\mathcal{O}}$ via $\sf{\mathcal{S}}$ are prevented by P4
Security Argument. We show that $\sf{SLAPP}$ satisfies both Definition IV and Definition IV implying adherence to Definition IV. Per Definition IV, IV-$\sf{PoX}$ security requires $\sf{PoX}$ assurance for stateless $\sf{\mathcal{F}}$ functions that run with input arguments, i.e., when ${\sf{\mathcal{S}}}=\phi$ and ${\sf{\mathcal{I}}}\neq\phi$. In this scenario, P2 and P5 ensure the atomic execution of $\sf{\mathcal{F}}$ with the authentic input $\sf{\mathcal{I}}$ provided by $\sf{\mathcal{V}rf}$, satisfying condition (i). P6 guarantees that $\sf{\mathcal{O}}$ is generated as a result of ${\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$ execution, fulfilling condition (ii). With both conditions met, $\sf{SLAPP}$ achieves IV-$\sf{PoX}$ security. Meanwhile, SV-$\sf{PoX}$ security in Definition IV requires secure $\sf{PoX}$ of a stateful $\sf{\mathcal{F}}$ executed without input arguments. Similar to the previous argument, P5 and P6 directly address conditions (i) and (ii) of SV-$\sf{PoX}$ security even when ${\sf{\mathcal{I}}}=\phi$. As P3 ensures $\sf{\mathcal{S}}$ privacy, it fulfills condition (iii). Also, P4 guarantees that $\sf{\mathcal{S}}$ cannot be modified except by a fresh instance of $\sf{SLAPP}$ protocol, which satisfies condition (iv). Meeting all conditions in Definition IV, $\sf{SLAPP}$ is also SP-$\sf{PoX}$-secure. Lastly, $\sf{PoSX}$ security (per Definition IV) follows directly from simultaneous adherence to IV-$\sf{PoX}$ and SP-$\sf{PoX}$ security.∎
VII $\sf{SLAPP}$ Extensions and Variations
Cryptographic Choices. Although Protocol V-B uses public-key cryptography, it can seamlessly transition to symmetric cryptography by simply substituting public-key operations (i.e., Sign and Verify) with MAC operations. If quantum threats are in scope, $\sf{SLAPP}$ can be similarly adjusted to support a post-quantum signature scheme. We demonstrate this versatility by implementing our prototype (Section IX) using three distinct cryptographic choices (public-key, symmetric, and post-quantum cryptography).
Space-Time Trade-Off. Section V-B discussed choices for managing ${\sf{\mathcal{S}}}_{sec}$ in the Secure World: (1) storing the entire $\sf{\mathcal{S}}$ value; or (2) maintaining a hash of $\sf{\mathcal{S}}$ value. The first prioritizes runtime efficiency since it requires no run-time hash computation while the second conserves storage in the Secure World by condensing the potentially large $\sf{\mathcal{S}}$ into a fixed-size digest. A third option, that eliminates storage in the Secure World, is to have SetState compute a MAC of $\sf{\mathcal{S}}$ (instead of a hash) and pass it back to the Non-Secure World for storage. Subsequently, the Secure World can authenticate $\sf{\mathcal{S}}$ received from the Normal World based on the MAC, yielding equivalent security guarantees as the previous two approaches. By default, $\sf{SLAPP}$ prototype adopts the second design choice, striking a balance between space and time overhead.
Multiple Stateful Functions. Our description of $\sf{SLAPP}$ assumes that $\sf{\mathcal{V}rf}$ intends to obtain one $\sf{PoSX}$ per $\sf{\mathcal{P}rv}$ at a time. Nonetheless, it can be extended to accommodate simultaneous $\sf{PoSX}$-s by maintaining a map between multiple ${\sf{\mathcal{S}}}_{sec}$-s and each ongoing $\sf{PoSX}$, in the Secure World. To that end, CheckState and SetState should be extended to assign ${\sf{\mathcal{S}}}_{sec}=map[id]$ for a given function identifier $id$. Also, in the last step of the protocol, $\sf{\mathcal{V}rf}$ must additionally inspect $\sf{\mathcal{F}}$ binary to ensure that it correctly calls CheckState and SetState with the correct function identifier.
Relaxing Atomicity Requirement. $\sf{SLAPP}$ security mandates atomicity (uninterruptability) during $\sf{\mathcal{F}}$ execution. This requirement may clash with real-time needs on $\sf{\mathcal{P}rv}$, potentially preventing time-sensitive tasks from completing while $\sf{SLAPP}$ is running. Recent studies [8, 40] propose techniques to relax this atomicity requirement in classic $\sf{PoX}$ and related schemes. These can also be adopted in $\sf{SLAPP}$ as follows: rather than completely disabling interrupts, the Secure World “locks” (i.e., making them read-only) $PMEM$, data currently in use by the $\sf{PoSX}$ task, and the Interrupt Vector Table (IVT). It also includes IVT in the snapshot $h$ before invoking $\sf{\mathcal{F}}$. As a consequence, the $\sf{PoSX}$ context is protected across interrupts. Once $\sf{\mathcal{F}}$ completes, respective memory can be unlocked.
VIII From $\sf{PoSX}$ to Poisoning Prevention
We now discuss how $\sf{SLAPP}$ can be leveraged to detect poisoning attacks. Unlike prior work, this guarantee extends to arbitrary functions $\sf{\mathcal{F}}$, whether they require private states or input arguments to execute. To illustrate this, we present ${}^{+}$ and ${}^{+}$, the poisoning-free versions of LDP-DC and FL-DC, respectively.
VIII-A Poisoning-free LDP.
${}^{+}$ utilizes $\sf{SLAPP}$ to ensure the correct execution of LDP phases:
  1. 1.
    Setup. Run $\sf{SLAPP}$ protocol to obtain a $\sf{PoSX}$ of Init-state that initializes $PRR$ mapping $B$ to an empty list on $\sf{\mathcal{P}rv}$. Abort if the protocol outputs $\bot$.
  2. 2.
    Collect. Run $\sf{SLAPP}$ protocol to obtain a $\sf{PoSX}$ of stateful LDP-DC function (in Algorithm 1) to perform a sensor reading and apply the LDP mechanism to it, returning the noisy output $\sf{\mathcal{O}}$ to $\sf{\mathcal{V}rf}$. Repeat this step if $\sf{\mathcal{V}rf}$ wants to collect more sensor readings.
$\sf{SLAPP}$ in Setup phase ensures that the state variable $B$ is initialized to an empty value. During Collect phase, $\sf{SLAPP}$ protocol gives assurance to $\sf{\mathcal{V}rf}$ that: (1) $\sf{\mathcal{O}}$ is genuine, originating from a timely execution of the sensor function, and correctly privatized by the underlying LDP mechanism, (2) $B$ corresponded to the authentic value (e.g., empty in the first time of this phase’s execution) right before and during the protocol execution, and (3) $B$ is correctly updated as a result of the protocol execution. These prevent poisoning attacks because $\sf{\mathcal{A}dv}$ can tamper with neither $\sf{\mathcal{O}}$ (during Collect) nor $B$ (during Setup or Collect).
VIII-B Poisoning-free FL.
${}^{+}$ follows a similar approach by using $\sf{SLAPP}$ to convey to $\sf{\mathcal{V}rf}$ the correct execution of:
  1. 1.
    Setup. Run $\sf{SLAPP}$ protocol to obtain a $\sf{PoSX}$ of Init-State function that initializes the local training dataset $D$ to an empty list on $\sf{\mathcal{P}rv}$. Abort if the protocol outputs $\bot$.
  2. 2.
    Collect. To perform a sensor reading and record the result to $D$, run $\sf{SLAPP}$ to obtain a $\sf{PoSX}$ of Sense-Store (see Algorithm 2). Repeat this step to collect more sensor readings. Abort if the protocol outputs $\bot$.
  3. 3.
    Train. Run $\sf{SLAPP}$ protocol to obtain a $\sf{PoSX}$ of Train (see Algorithm 2) that performs a local training on $D$ and output the locally trained model $\sf{\mathcal{O}}$ to $\sf{\mathcal{V}rf}$.
Similar to ${}^{+}$, Setup guarantees to $\sf{\mathcal{V}rf}$ that $D$ starts empty. For Collect, $\sf{SLAPP}$ ensures that each record in $D$ is produced as a result of executing the expected sensor function on $\sf{\mathcal{P}rv}$. Finally, Train assures $\sf{\mathcal{V}rf}$ that the correct training function was applied to the authentic training data in $D$, leading to the received trained model $\sf{\mathcal{O}}$. In the context of FL, $\sf{\mathcal{A}dv}$ may attempt poisoning attacks in two ways: (1) model poisoning by directly tampering with the trained model $\sf{\mathcal{O}}$ and (2) data poisoning by manipulating the dataset $D$ used for training. Both are prevented by ${}^{+}$.
Remark. We emphasize that both ${}^{+}$ and ${}^{+}$ maintain the privacy of their original counterparts. The only additional information obtained by $\sf{\mathcal{V}rf}$ is $\sigma$, which is not a function of the underlying execution state. Also, as a system-level approach, these schemes offer a deterministic guarantee in differentiating poisoned data from authentic data, i.e., achieving 100% true positive and true negative rates, without relying on any assumptions about the underlying data.
IX Evaluation
IX-A Experimental Setup
Cryptographic Variants: As noted in Section VII, $\sf{SLAPP}$ is agnostic to underlying cryptographic primitives. To showcase this flexibility, we provide 3 implementation variants of $\sf{SLAPP}$’s RoT: $\sf{SLAPP}_{SK}$, $\sf{SLAPP}_{PK}$, and $\sf{SLAPP}_{PQ}$:
  1. 1.
    $\sf{SLAPP}_{SK}$ uses symmetric-key cryptography to implement Sign and Verify in Protocol V-B with an HMAC-SHA256 in line with prior $\sf{PoX}$ work [13].
  2. 2.
    $\sf{SLAPP}_{PK}$ relies on the public-key signature ECDSA NIST256p from micro-ECC library [^1] https://github.com/kmackay/micro-ecc, which is commonly used in embedded settings [49, 39].
  3. 3.
    $\sf{SLAPP}_{PQ}$ implements Sign and Verify using the quantum-resistant public-key signature Sphincs+, a low-RAM version of Sphincs-sha2-128f [^2] https://github.com/sphincs/low-ram-sphincsplus. This choice is supported by prior research [41, 26] showing feasibility of Sphincs+ on the target ARM Cortex-M devices.
Prototype. We prototype $\sf{SLAPP}$ variants on SAML11 Xplained Pro [36] development board as a representative of resource-constrained IoT devices. It features TrustZone-M on an ARM Cortex-M23 MCU @ 32MHz clock, with 64KB of FLASH (of which we assign 32KB to store the Non-secure World’s PMEM) and 16KB of RAM. To accurately isolate $\sf{SLAPP}$ base overhead, we implement a simple stateful $\sf{\mathcal{F}}$ function on $\sf{\mathcal{P}rv}$ that takes input from a GPIO port specified from a $\sf{PoSX}$ request, performs a sensor reading on that port, and outputs an accumulative sum of all readings over time to $\sf{\mathcal{V}rf}$. $\sf{\mathcal{V}rf}$ is deployed as a commodity desktop equipped with an Intel i7-8809G CPU @ 3.10GHz and 32GB RAM. $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$ are connected via serial communication. All prototypes are open-sourced and publicly available at [3].
TABLE I: Lines of code and binary size of $\sf{SLAPP}$’s TCB. Overhead w.r.t. baseline is shown in parentheses.
Variants LoC Bin. Size in KB
$\sf{SLAPP}_{SK}$ 605 (+5.6%) 4.62 (+5.8%)
$\sf{SLAPP}_{PK}$ 1490 (+2.28%) 13.06 (+2.0%)
$\sf{SLAPP}_{PQ}$ 2806 (+1.24%) 13.28 (+2.1%)
IX-B Baseline.
For comparison, we consider an alternative naive baseline approach in which all functions to prove execution $\sf{\mathcal{F}}$ are included as part of TCB residing within the Secure World. To perform $\sf{PoSX}$, TCB receives and authenticates a $\sf{PoSX}$ request from the Non-Secure World, just like $\sf{SLAPP}$. Unlike $\sf{SLAPP}$, this approach executes $\sf{\mathcal{F}}$ inside the Secure World, The result is signed (or MAC’ed) and forwarded to $\sf{\mathcal{V}rf}$, akin to $\sf{SLAPP}$. As this baseline approach is also agnostic to the underlying cryptographic primitives, we refer to $\sf{Baseline_{SK}}$, $\sf{Baseline_{PK}}$ and $\sf{Baseline_{PQ}}$ as the baseline approaches that utilize symmetric-key cryptography (HMAC-SHA256), public-key cryptography (ECDSA) and post-quantum cryptography (Sphincs+), respectively.
We note that this baseline faces several security and practical downsides. First, its TCB becomes bloated and dependent on multiple untrusted applications by including all $\sf{\mathcal{F}}$-s within the Secure World. This implies that a vulnerability in one of them can compromise all (i.e., violating the principle of least privilege). It also incurs Secure World’s additional storage for maintaining data of $\sf{\mathcal{F}}$ execution. This reduces available RAM for normal applications in the Non-Secure World. Moreover, it makes the Secure World code (which should be immutable post-deployment – recall Section II-B) application-specific and thus requires rewriting/updating the Secure World every time to support new applications, which may necessitate cumbersome physical intervention.
IX-C Space Overhead
Code Size. $\sf{SLAPP}$’s TCB was implemented in C and compiled using the O3 optimization flag. Details of code size and the resulting binaries are shown in Table I. As $\sf{SLAPP}_{SK}$ relies on an inexpensive cryptographic primitive, it exhibits the smallest code size. Conversely, $\sf{SLAPP}_{PQ}$ results in the largest number of lines of C code. However, the O3 optimization was capable of reducing its binary size, making it comparable to that of $\sf{SLAPP}_{PK}$. Compared to the baseline, $\sf{SLAPP}$ introduces a small code size overhead, approximately 6% for $\sf{SLAPP}_{SK}$ and 2% for the others.
For $\sf{\mathcal{F}}$ instrumentation, $\sf{SLAPP}$ requires prepending a call to CheckState at the beginning of $\sf{\mathcal{F}}$ and another call to SetState before $\sf{\mathcal{F}}$ returns. This instrumentation results in only 2 additional lines of C code, enlarging $\sf{\mathcal{F}}$’s binary (residing in Non-Secure World) by a fixed 44 bytes.
Figure 4: Peak runtime data allocation across all approaches
Memory Usage. Next, we measure the peak amount of Secure World data allocated at runtime. This data consists of all stack and static/global variables (our implementation utilizes no heap allocation). As shown in Figure 4, $\sf{SLAPP}_{SK}$ and $\sf{SLAPP}_{PK}$ use roughly the same amount of data: $\approx$1.5KB. $\sf{SLAPP}_{PQ}$, on the other hand, requires more than double this amount: $\approx$3.6KB. Compared to the baseline, our $\sf{SLAPP}$ variants incur an additional memory usage ranging from 32 bytes for $\sf{SLAPP}_{PQ}$ to 200 bytes for $\sf{SLAPP}_{SK}$. These correspond to $<$ 1.25% of available RAM.
Figure 5: Execution time across all approaches
IX-D Time Overhead
As $\sf{\mathcal{V}rf}$ operates on a powerful back-end, its runtime within the $\sf{SLAPP}$ protocol is negligible compared to the execution time on $\sf{\mathcal{P}rv}$. Thus, we focus on measuring $\sf{\mathcal{P}rv}$’s execution time (i.e., the runtime time of executing Phases 3, 4, and 5 in Section V). Results are illustrated in Figure 5.
As expected, $\sf{SLAPP}_{SK}$ has the fastest runtime: less than half a second. On the other hand, $\sf{SLAPP}_{PQ}$ utilizes expensive cryptography and thus incurs the longest execution time of $\approx 100$ seconds. $\sf{SLAPP}_{PK}$ positions between these two, taking around 2.5 seconds.
Figure 6: Execution time with varying $PMEM$ size
Since $\sf{SLAPP}$ requires $\sf{\mathcal{P}rv}$ to compute a snapshot of $PMEM$ as part of proof generation, we conducted an experiment to assess the impact of $PMEM$ size on $\sf{\mathcal{P}rv}$’s execution time. In this experiment, we varied the $PMEM$ size from 4KB to 32KB. As depicted in Figure 6, the time to execute this snapshot contributes significantly to the overall execution time when $\sf{SLAPP}_{SK}$ is used. It also shows a linear relationship between execution time and $PMEM$ size for $\sf{SLAPP}_{SK}$. However, for $\sf{SLAPP}_{PK}$ and $\sf{SLAPP}_{PQ}$, this effect is negligible; its runtime has minimal impact on the overall execution time regardless of $PMEM$ size.
Also, from Figures 5 and 6, we can conclude that all $\sf{SLAPP}$ variants incur a negligible execution time overhead, i.e., $<1\%$, compared to their baseline counterparts.
IX-E Case Study 1: Local Differential Privacy
In the first case study, we envision the integration of $\sf{SLAPP}$ within a smart city/grid system. The service provider aims to periodically collect energy consumption data from all smart meters located in individual households to calculate electric bills, forecast load, etc. Previous studies [35, 38] have demonstrated privacy risks by exposing raw energy to the service provider, e.g., with access to such data, the provider could potentially infer users’ habits and behaviors.
To address this privacy concern, one can employ LDP-DC from Algorithm 1. The service provider is also motivated to employ a poisoning-free version, ${}^{+}$ as presented in Section VIII-A, to prevent poisoning by detecting malicious outputs from edge devices.
TABLE II: Performance of FL/LDP with $\sf{SLAPP}_{SK}$
Protocol Task Phase Runtime Non-Secure Secure
Mem. Usage Mem. Usage
${}^{+}$ Collect noisy Setup 0.50s 1KB 1.48KB
energy readings Collect 2.30s 1.27KB 1.48KB
${}^{+}$ Train a classifier for detecting anomalous energy usage patterns Setup 0.50s 2.57KB 1.48KB
Collect 1.98s 2.7KB 1.48KB
Train 21.84s 3.97KB 1.48KB
We build a prototype of a smart meter ($\sf{\mathcal{P}rv}$) based on a SAML11 Xplained Pro development board connecting to a PZEM-004T energy meter module. $\sf{\mathcal{P}rv}$’s Non-Secure World consists of LDP-DC software (Algorithm 1) and a driver responsible for retrieving energy data from the PZEM-004T module. Meanwhile, the Secure World contains $\sf{SLAPP}_{SK}$’s TCB. The service provider ($\sf{\mathcal{V}rf}$) runs on a commodity desktop and communicates with $\sf{\mathcal{P}rv}$ over serial communication.
During normal (benign) operation, $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$ execute an instance of ${}^{+}$ protocol. This instance begins with the Setup phase, which requires $\sf{PoSX}$ of Init-State function to initialize a $\sf{\mathcal{P}rv}$-local state ($PRR$ mapping) to zeroes. Since $\sf{\mathcal{P}rv}$ faithfully executes this phase, it completes with $\top$. Upon obtaining the successful output, $\sf{\mathcal{V}rf}$ proceeds to the Collect phase, which aims to collect authentic noisy energy data from $\sf{\mathcal{P}rv}$. To achieve this, $\sf{\mathcal{V}rf}$ crafts a $\sf{PoSX}$ request by configuring a function to prove execution to LDP-DC and properly selecting LDP input parameters (i.e., $f$, $p$ and $q$) to meet the $\epsilon$-LDP requirement, and sends an authenticated $\sf{PoSX}$ request to $\sf{\mathcal{P}rv}$. Adhering to the protocol, $\sf{\mathcal{P}rv}$ executes LDP-DC correctly (with $\sf{\mathcal{V}rf}$-defined inputs) and returns authentic noisy energy data to $\sf{\mathcal{V}rf}$. As a result, this phase completes with $\top$ ensuring $\sf{\mathcal{V}rf}$ that the received energy data is not poisoned. End-to-end runtime and memory usage during this operation are shown in Table II.
Next, we launch the following attacks to ${}^{+}$:
$\bullet$ ${\sf{\mathcal{A}dv}}_{1}$ corrupts Init-State to poison the initial state.
$\bullet$ ${\sf{\mathcal{A}dv}}_{2}$ corrupts LDP-DC between Setup and Collect phases to poison raw energy readings.
$\bullet$ ${\sf{\mathcal{A}dv}}_{3}$ poisons the state in between Setup and Collect phases.
$\bullet$ ${\sf{\mathcal{A}dv}}_{4}$ poisons the final result in Collect phase.
${}^{+}$ detects all aforementioned attacks. Specifically, ${\sf{\mathcal{A}dv}}_{1}$ and ${\sf{\mathcal{A}dv}}_{2}$ are detected at the end of Setup and Collect phases, respectively, due to the software binary mismatch. ${\sf{\mathcal{A}dv}}_{3}$ is caught in Collect since $\sf{SLAPP}$’s RoT noticed tampered state values (from CheckState) and thus refused to generate a valid proof. Similarly, ${\sf{\mathcal{A}dv}}_{4}$ is detected in Collect phase since the proof $\sigma$ does not reflect the tampered output.
IX-F Case Study 2: Federated Learning
Here the service provider aims to train an ML model for identifying anomalous usage patterns in a smart grid system. To achieve this while ensuring user privacy, it employs FL based on a neural network, as described in [27]. Each smart meter independently trains a local neural network model with 8 hidden units on its local data. These local models are later transmitted to and aggregated by the service provider. ${}^{+}$ is employed to prevent poisoning.
The setup resembles the previous case study, where a smart meter device ($\sf{\mathcal{P}rv}$) operates on a SAML11 Xplained Pro board. It performs energy readings by interfacing with a PZEM-004T device and connects to $\sf{\mathcal{V}rf}$ via serial communication.
During benign operation, $\sf{\mathcal{P}rv}$ and $\sf{\mathcal{V}rf}$ initiate ${}^{+}$ protocol by running Setup phase, which clears all sensor readings on $\sf{\mathcal{P}rv}$. Next, $\sf{\mathcal{V}rf}$ periodically triggers Collect phase to record an energy reading into a $\sf{\mathcal{P}rv}$-local training dataset. Once $\sf{\mathcal{P}rv}$ records enough energy readings, $\sf{\mathcal{V}rf}$ triggers the Train phase by transmitting a $\sf{PoSX}$ request to $\sf{\mathcal{P}rv}$. This request specifies initial weights to be trained on, a learning rate, and the number of epochs to be 100. $\sf{\mathcal{P}rv}$ then performs the local training. Upon completion, $\sf{\mathcal{P}rv}$ sends back the updated weights along with the proof of training to $\sf{\mathcal{V}rf}$. Since $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$ adhere to ${}^{+}$, the protocol yields $\top$, ensuring to $\sf{\mathcal{V}rf}$ that the received model was correctly trained on authentic energy data and can thus be safely aggregated onto the global model. Runtime breakdown and memory usage are shown in Table II.
We simulate two types of adversaries: ${\sf{\mathcal{A}dv}}_{D}$ and ${\sf{\mathcal{A}dv}}_{M}$. ${\sf{\mathcal{A}dv}}_{D}$ performs data poisoning attacks to ${}^{+}$ by compromising $\sf{\mathcal{P}rv}$-local training dataset before Train is executed. Meanwhile, ${\sf{\mathcal{A}dv}}_{M}$ simulates model poisoning by manipulating the locally trained model parameters before the model arrives at $\sf{\mathcal{V}rf}$. ${}^{+}$ detects ${\sf{\mathcal{A}dv}}_{D}$ during the Train phase because $\sf{SLAPP}$’s RoT identifies state $\sf{\mathcal{S}}$ tampering ($\sf{\mathcal{S}}$ is the training dataset in this case). ${\sf{\mathcal{A}dv}}_{M}$ is caught at the end of the Train as the proof does not match the tampered output.
X Related Work
Poisoning prevention LDP and FL. Aside from data-driven  [57, 23, 6]) and algorithmic [50, 32, 55, 33] poisoning prevention approaches discussed in Section I, a few recent schemes propose system security approaches [11, 48, 2]. To the best of our knowledge, none of these schemes provide $\sf{PoSX}$-equivalent guarantee of provable integrity all the way from data acquisition until its de facto usage as an aggregated statistical result or global ML model. They also do not address the issues of input validation or state preservation for arbitrary functions $\sf{\mathcal{F}}$ to be execute on $\sf{\mathcal{P}rv}$.
Verifiable Software Integrity in Embedded Devices. To secure low-end embedded devices, various low-cost security architectures have been proposed for remote verification of their software state via integrity proofs [44]. These proofs vary in terms of expressiveness, with simpler ones confirming correct binary presence (remote attestation) [12, 21, 5, 29, 20, 42], while more expressive ones support verification of arbitrary code execution. Aside from $\sf{PoX}$ architectures  [13, 8, 45, 34], control flow attestation/auditing (CFA) techniques  [40, 1, 52, 60, 51, 56, 18, 17, 59, 14, 9] prove to $\sf{\mathcal{V}rf}$ the exact order in which instructions have executed within a particular code in $\sf{\mathcal{P}rv}$, thus enabling detection of code reuse attacks that can be triggered if the code whose execution is being proven is itself vulnerable. Data flow attestation [51, 15, 17] augments CFA to generate evidence about memory safety violations even when exploits do not alter a program’s legal control flow path.
Private Data Collection on Edge Devices. Complementary to privacy mechanisms focusing on hiding private data from back-ends (e.g., LDP/FL), recent work has delved into assuring that private data is secure against compromised sensing devices “from its birth”, i.e., from the moment when it is digitized. VERSA [43] was proposed as a HW/SW architecture to guarantee that only the correct execution of expected and explicitly authorized software can access and manipulate sensing interfaces. As a consequence, it blocks malware/modified software from accessing sensitive sensed quantities by default. Following this notion, Sensing And Actuation As A Privilege (SA4P) [16] realizes this concept using ARM TrustZone.
XI Conclusion
We defined and developed stateful proofs of execution, a system security primitive to thwart poisoning in applications such as differential privacy and federated learning. We analyze the security of our design ($\sf{SLAPP}$) and evaluate its performance with an open-source prototype. Results indicate strong poisoning prevention guarantees at modest overhead applicable even to MCU-based resource-constrained IoT devices.